# Show-o: One Single Transformer to Unify Multimodal Understanding and Generation

### v6

@@ 5.5 Ablation Studies
The default Show-o employs MAGVIT-v2 to encode images into discrete tokens for both multimodal understanding and generation. Inspired by the literature (Liu et al., 2024b), we investigate the impact of the most popular design choice of vision encoder, i.e., the pre-trained CLIP ViT (Radford et al., 2021), for multimodal understanding. We first compare the two settings using our Show-o model. In Table 5, the comparison between Exp 2 and Exp 4, Exp 3 and Exp 5 clearly demonstrates that continuous representations from CLIP-ViT have significantly better performance on multimodal understanding than that of MAGVIT-v2. This mainly attributes to: i) The CLIP-ViT is pre-trained on a much larger dataset (400M) than that of our pre-trained MAGVIT-v2 (35M); ii) In contrast to image reconstruction learning objective in MAGVIT-v2, the discriminative loss, i.e., image-text matching, in CLIP-ViT makes the extracted representations easier to be adapted for multimodal understanding.
# Image-text | Image Resolution | # Image Tokens | POPE | MME | Flickr30k | VQAv2 | GQA | MMMU |
35M | 2562 | 256 | 73.8 | 948.4 | 36.2 | 59.3 | 48.7 | 25.1 |
2.0B | 2562 | 256 | 76.2 | 1014.9 | 48.9 | 64.7 | 54.2 | 25.0 |
2.0B | 5122 | 1024 | 80.0 | 1097.2 | 62.5 | 69.4 | 58.0 | 26.7 |
@@
1 | LLaVA | CLIP-ViT | ✗ | Continuous | 84.1 | 1128.0 | 69.6 | 73.0 | 56.5 | 30.67 |
2 | Show-o‡ | CLIP-ViT | ✓ | Continuous | 84.5 | 1182.7 | 64.3 | 71.9 | 57.5 | 27.4 |
3 | Show-o‡ | CLIP-ViT | ✗ | Continuous | 84.5 | 1161.6 | 68.5 | 73.5 | 58.7 | 29.2 |
4 | Show-o† | MAGVIT-v2 | ✓ | Continuous | 74.3 | 947.8 | 33.9 | 59.4 | 51.0 | 26.7 |
5 | Show-o† | MAGVIT-v2 | ✗ | Continuous | 65.1 | 800.0 | 12.3 | 50.8 | 43.9 | 24.6 |
6 | Show-o | MAGVIT-v2 | ✓ | Discrete | 73.8 | 948.4 | 36.2 | 57.8 | 48.7 | 25.1 |
7 | Show-o | MAGVIT-v2 | ✗ | Discrete | 63.8 | 689.1 | 4.5 | 46.1 | 40.5 | 28.1 |
Impact of Various Representations for Multimodal Understanding.
In typical multimodal understanding models like LLaVA, the image representation extraction and cross-modal alignment usually happen in the continuous space. However, image tokenizers such as MAGVIT-v2 naturally yield discrete image tokens. As shown in Table 5, we compare the two types of input, i.e., continuous representations and discrete tokens, in the multimodal understanding scenario. In Exp 6 and 7, we use the pre-trained MAGVIT-v2 to extract discrete tokens and train an embedding layer to embed the tokens into the continuous embedding space of the LLM. In Exp 4 and 5, we modify MAGVIT-v2 to output continuous representations without quantization. The cross-modal projection layer follows the setting of LLaVA. The comparison between Exp 5 and Exp 7 reveals that discrete tokens show much worse performance on most benchmarks. We attribute the performance gap to that popular multimodal understanding datasets, e.g., LLaVA-Pretrain-558K, are not sufficient to align discrete image tokens into the language space, leading to an unsatisfactory cross-modal understanding. In contrast, continuous representations, already lying in a well-shaped embedding space, are much easier to align.
Impact of Unified Pre-training for Multimodal Understanding.
Our training pipeline involves two-stage unified pre-training to learn image token embedding and image-text alignment for multimodal understanding and generation (as described in Section 4.3). Here we elaborate on the impact of the unified per-training with different vision encoders and types of representations:
@@
Additionally, we present qualitative examples to illustrate the impact of sampling steps and classifier-free guidance for text-to-image generation in the following.
@@
5.6 Failure Cases
@@
6 Conclusion
7 Acknowledgments
We would like to express our sincere gratitude to Henry Hengyuan Zhao for his valuable discussions and insightful feedback on multimodal understanding and Mingrui Wang for his patient assistance in helping us set up the development environment.
Appendix A Appendix
A.1 Alternative Lower Bound for the variational diffusion
The constants and are:
The alternative lower bound can be derived as:
This proof is provided by Zhang et al. (2024).
A.2 Image Generation details
Given text tokens and tokens as initial input, Show-o predict logits in parallel, where and is the time step. For each token at location , we sample an image token from the codebook based on the predicted probability and indicate its predicted score as the confidence of this sampled token. The confidence of the unmasked token is set as 1.0. Next, we compute the number of image tokens that should be re-masked based on the mask scheduling function , where . More details of mask scheduling functions can be found in MaskGIT (Chang et al., 2022). Subsequently, we replace the predicted image tokens with token based on the following metric:

### v7

@@ Appendix I Ablation Studies
Impact of Vision Encoder for Multimodal Understanding. The default Show-o employs MAGVIT-v2 to encode images into discrete tokens for both multimodal understanding and generation. Inspired by the literature (Liu et al., 2024b), we investigate the impact of the most popular design choice of vision encoder, i.e., the pre-trained CLIP ViT (Radford et al., 2021), for multimodal understanding. We first compare the two settings using our Show-o model. In Table 5, the comparison between Exp 2 and Exp 4, Exp 3 and Exp 5 clearly demonstrates that continuous representations from CLIP-ViT have significantly better performance on multimodal understanding than that of MAGVIT-v2. This mainly attributes to: i) The CLIP-ViT is pre-trained on a much larger dataset (400M) than that of our pre-trained MAGVIT-v2 (35M); ii) In contrast to image reconstruction learning objective in MAGVIT-v2, the discriminative loss, i.e., image-text matching, in CLIP-ViT makes the extracted representations easier to be adapted for multimodal understanding.
@@
| 1 | LLaVA | CLIP-ViT | ✗ | Continuous | 84.1 | 1128.0 | 69.6 | 73.0 | 56.5 | 30.67 |
| 2 | Show-o‡ | CLIP-ViT | ✓ | Continuous | 84.5 | 1182.7 | 64.3 | 71.9 | 57.5 | 27.4 |
| 3 | Show-o‡ | CLIP-ViT | ✗ | Continuous | 84.5 | 1161.6 | 68.5 | 73.5 | 58.7 | 29.2 |
| 4 | Show-o† | MAGVIT-v2 | ✓ | Continuous | 74.3 | 947.8 | 33.9 | 59.4 | 51.0 | 26.7 |
| 5 | Show-o† | MAGVIT-v2 | ✗ | Continuous | 65.1 | 800.0 | 12.3 | 50.8 | 43.9 | 24.6 |
| 6 | Show-o | MAGVIT-v2 | ✓ | Discrete | 73.8 | 948.4 | 36.2 | 57.8 | 48.7 | 25.1 |
| 7 | Show-o | MAGVIT-v2 | ✗ | Discrete | 63.8 | 689.1 | 4.5 | 46.1 | 40.5 | 28.1 |
Impact of Various Representations for Multimodal Understanding. In typical multimodal understanding models like LLaVA, the image representation extraction and cross-modal alignment usually happen in the continuous space. However, image tokenizers such as MAGVIT-v2 naturally yield discrete image tokens. As shown in Table 5, we compare the two types of input, i.e., continuous representations and discrete tokens, in the multimodal understanding scenario. In Exp 6 and 7, we use the pre-trained MAGVIT-v2 to extract discrete tokens and train an embedding layer to embed the tokens into the continuous embedding space of the LLM. In Exp 4 and 5, we modify MAGVIT-v2 to output continuous representations without quantization. The cross-modal projection layer follows the setting of LLaVA. The comparison between Exp 5 and Exp 7 reveals that discrete tokens show much worse performance on most benchmarks. We attribute the performance gap to that popular multimodal understanding datasets, e.g., LLaVA-Pretrain-558K, are not sufficient to align discrete image tokens into the language space, leading to an unsatisfactory cross-modal understanding. In contrast, continuous representations, already lying in a well-shaped embedding space, are much easier to align.
Impact of Unified Pre-training for Multimodal Understanding. Our training pipeline involves two-stage unified pre-training to learn image token embedding and image-text alignment for multimodal understanding and generation (as described in Section C). Here we elaborate on the impact of the unified per-training with different vision encoders and types of representations:
@@
Appendix J Mask-free image editing
Given an image, it can be converted into discrete image tokens, which can then be subject to iterative random masking for sampling purposes. This process allows for image editing without the need for predefined masks. For instance, as illustrated in Fig. 13, Show-o enables local-region modifications such as ”changing the red apple to green” or ”replacing the apple with a mouse.” Furthermore, Show-o facilitates global style adjustments like ”transforming the original image into a cartoon or oil painting style.”
Appendix K Failure Cases